Animation Source Forums

=> Visit the Animation Source fansites!

Read the site rules!

All times are UTC - 6 hours




Post new topic Reply to topic  [ 1 post ] 
Author Message
 Post subject: Difference Between Probability and Odds
PostPosted: Tue Aug 02, 2022 9:04 am 

Joined: Sat Dec 14, 2019 11:08 am






Probability and odds are very closely related; the terms are often used interchangeably, but there is an important difference.

Probability: divide chances of winning by the total number of chances available . For example, if you buy one ticket for a raffle with 100 tickets sold, you have one possible chance at a win, with 100 possible chances overall. Your probability of winning is 1/100. Probabilities can also be expressed as:
Decimals; 1/100 = 0.01 (a fraction is just a division problem: plug 1/100 into any calculator to get the decimal equivalent)
Percentages: to convert to a percent, multiply your decimal by 100 and add a % sign: 0.01 * 100 = 1%.
Odds are a ratio of your chances of losing to your chances of winning. Using the above raffle example, your chances of losing are 99 (the “other” tickets) and your chances of winning are 1 (the ticket you purchased). Your odds are 99 to 1.
To convert odds to probability:

Place your chance of winning (1 in this example) in the numerator (top) of a fraction.
1/?
Add both values in the odds (99 and 1 in this example) and place that value in the fraction’s denominator (bottom): 1 + 99 = 100.
1/100
How To Calculate Odds of Winning: Example
You’ll often find sweepstakes listing odds in their official rules, but they are normally quite high. Sometimes, the odds will be explicitly stated. For example, Quilted Northern Bathroom Tissue’s 2005 sweepstakes published odds of 1:11,000,000 [1]. But in the vast majority of cases, you’ll find this statement in the official rules, like this one from Ellen’s 2021 sweepstakes [2]:

“Odds of winning depend upon the number of eligible entries received during the Entry Period.”

In order to calculate odds of winning (or probability of winning), you have to know how many total entries. Thanks to the advent of the internet, these contests gather millions of entries, which means your odds are extremely small of winning a particular sweepstakes. Let’s take a look at an example.

HGTV’s 2021 Smart Home sweepstakes, which offered the winner $100,000, a 2021 Mercedes-Benz, and a brand-new smart home in Naples, Florida, had 106 million entries [3]. You can use that information to calculate your odds of winning. The contest, which runs every year, allows each person two entries a day for around six weeks (42 days).


Top
 Profile  
Reply with quote  
Display posts from previous:  Sort by  
Post new topic Reply to topic  [ 1 post ] 

All times are UTC - 6 hours



You cannot post new topics in this forum
You cannot reply to topics in this forum
You cannot edit your posts in this forum
You cannot delete your posts in this forum
You cannot post attachments in this forum

Search for:
Jump to:  
Powered by phpBB © 2000, 2002, 2005, 2008 phpBB Group